OpenStudy (anonymous):

In how many ways can 2 different ice cream toppings be chosen from 12 available toppings?

OpenStudy (ash2326):

@anlnt4 Do you know combinations?

OpenStudy (anonymous):

we're learning them now, I think. I know the answer, I just don't know how to get it.

OpenStudy (anonymous):

What is the answer?

OpenStudy (anonymous):

66

OpenStudy (anonymous):

12 choices for the first topping, then 11 for the second topping so \(12\times 11\) exept that the the order you choose does not matter, so divide that by 2

OpenStudy (anonymous):

in other words it is \(\frac{12\times 11}{2}=6\times 11=66\)

OpenStudy (anonymous):

So anytime the order doesn't matter you'd divide it by the number you need?

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):